### v1

# Unknotting number is not additive under connected sum

Abstract.
We give the first examples of a pair of knots $K_{1}$,$K_{2}$ in the 3-sphere for which their unknotting numbers satisfy $u(K_{1}\#K_{2})<u(K_{1})+u(K_{2})$ This answers question 1.69(B) from Kirby’s problem list [16], in the negative.
Mathematics Subject Classification 2020: 57K10, 57K31
1. Introduction
The unknotting number (or Gordian number) $u(K)$ of a knot $K$ is one of the most fundamental measures of the complexity of a knot. It is defined as the minimum number of crossing changes, interspersed with isotopy, required to transform a diagram of $K$ to a diagram of the unknot. A wide array of techniques have been employed over the years to compute unknotting numbers, using nearly every new technique that has been introduced into knot theory (see [8], [15], [17], [18], [20], [21], [22], [23], [25], [26] for a selection). Yet to this day there are still nine prime 10-crossing knots whose unknotting numbers remain unknown, a testament to the difficulty of determining this basic invariant. See the Knotinfo site [19] for the most up-to-date list of unresolved knots.
Connected sum is a basic operation for creating or decomposing knots and links. A diagram for the connected sum $K_{1}\#K_{2}$ of two knots $K_{1}$ and $K_{2}$ can be obtained from disjoint diagrams of the two knots by deleting a short arc in each and introducing two new arcs to connect the endpoints of the deleted arcs to create a knot, without introducing additional crossings in the diagram. A knot is prime if it cannot be expressed as the connected sum of two non-trivial knots. Every knot admits a unique decomposition (up to ordering) as a connected sum of prime knots [31]. This has enabled many problems in knot theory to focus on prime knots as their principle object of study.
Many knot-theoretic constructions and invariants are known to behave well under the connected sum operation. Polynomial invariants, such as the Alexander [1], Jones [14], and HOMFLY polynomials [12] are multiplicative under connected sum, and many basic numerical invariants are very well-behaved under conneccted sum, such as the 3-genus [29], signature [21], braid index [4], and bridge number [30]. Unknotting number has long been conjectured to be additive under connected sum; this conjecture is implicit in the work of Wendt, in one of the first systematic studies of unknotting number [37]. It is unclear when and where this was first explicitly stated; most references to it call it an ‘old conjecture’. It can be found in the problem list of Gordon [13] from 1977 and in Kirby’s list [16].
Conjecture 1.1.
For knots $K_{1}$ and $K_{2}$ in the 3-sphere, $u(K_{1}\#K_{2})=u(K_{1})+u(K_{2})$.
Showing that $u(K_{1}\#K_{2})\leq u(K_{1})+u(K_{2})$ is very straighforward; choose diagrams for $K_{1}$ and $K_{2}$ that can be unknotted with $u(K_{1})$ and $u(K_{2})$ crossing changes, respectively, and create the connected sum $K_{1}\#K_{2}$ using these diagrams. Then the same set of $u(K_{1})+u(K_{2})$ crossing changes will transform $K_{1}\#K_{2}$ into (unknot)##\##(unknot) = the unknot. The challenge has always been, essentially, to determine whether one cannot do better than this approach, and so $u(K_{1}\#K_{2})\geq u(K_{1})+u(K_{2})$, as well.
Progress on this conjecture has been slow, however. Perhaps the most significant work in support of it has been Scharlemann’s result [28] that a knot $K$ with $u(K)=1$ is prime; consequently, the connected sum of two non-trivial knots must have unknotting number at least 2. More recently, Yang [38] has shown that the connected sum of $n$ identical knots $K$ whose Alexander polynomial satisfies $\Delta_{K}(t)\neq 1$ has unknotting number at least $n$. Very recently, Applebaum, Blackwell, Davies, Edlich, Juhász, Lackenby, Tomašev, and Zheng [2] employed reinforcment learning to search for counterexamples to this conjecture, and established some interesting consequences for unknotting sequences of connected sums. They also showed how, if the conjecture above were true, one could compute the unknotting numbers of many prime knots.
It was this last perspective - what can we learn if we assume that unknotting number is additive - which inpired the authors of the present work to carry out the computations that led to their discovery. We use the notation $\overline{K}$ to represent the mirror image of the knot $K$. It is an elementary fact that $u(L)=u(\overline{L})$ for every knot $L$.
Theorem 1.2.
Unknotting number is not additive under connected sum. In particular, for the knot $K=7_{1}$ (the (2,7)-torus knot), with $u(K)=3$, the connected sum $L=K\#\overline{K}$ satifies $u(L)\leq 5<6=u(K)+u(\overline{K})$.
The additivity of unknotting number appears as Problem 1.69(B) in Kirby’s problem list [16]; this result therefore settles this question (in the negative).
Once we have this one example, it takes little work to find infinitely many more. Because the torus knot $7_{1}=T(2,7)$ appears in a minimal unknotting sequence for each of the torus knots $T(2,2k+1)$, $k\geq 3$, for example, we can show:
Corollary 1.3.
For every natural number $k,\ell\geq 3$,
$u(T(2,2k+1)\#\overline{T(2,2\ell+1)})<u(T(2,2k+1))+u(T(2,2\ell+1))\ .$
See Section 2 for more examples.
Our approach to discovering our main example, described in section 2 below, is modeled on the authors’ previous work settling the Bernhard-Jablan Conjecture [6], and at least partly relies on data generated from that project. We also must credit, and express our immense gratitude to, the authors and developers of Snappy/SnapPea [9]; no part of this project would have been possible without the extensive computations made possible by this software.
The structure of this paper is as follows. In Section 2 we discuss the knots $K=7_{1}$ and $L=K\#\overline{K}$, and give the knot diagrams and crossing changes which establish the failure of additivity for unknotting number. In Section 3 we describe the motivation for our searches and the process by which we carried them out. In Section 4 we discuss the consequences of this example for the computation of unknotting number, and possible future directions. The final section provides python code which will allow the reader to carry out their own verification of the computations used in our main example.
2. Proofs of Theorem 1.2 and Corollary 1.3
The two summands which make up our example in Theorem 1.2 are $K=7_{1}$ and its mirror image $\overline{K}$. $K$ is the $(2,7)$-torus knot. A routine calculation establishes that $K$ has signature $\sigma(K)=6$; a result of Murasugi [21] gives $|\sigma(K)|/2\leq u(K)$, and so $K$ has unknotting number at least 3. Since any 3 crossing changes in its standard 7-crossing diagram results in the unknot, we have $u(K)=3$. Alternatively, one can appeal to the important and more far-reaching result of Kronheimer and Mrowka [17] that the unknotting number of the $(p,q)$-torus knot $T(p,q)$ is $(p-1)(q-1)/2$. It is an elementary result that unknotting number is preserved under mirror image - hold a mirror up to your unknotting sequences - and so $u(\overline{K})=3$, as well. Consequently, $u(K)+u(\overline{K})=3+3=6$.
We will demonstrate, however, that $L$ has unknotting number at most $5$. We do this by giving a diagram of $L$ for which changing two crossings gives (a diagram for) the knot $K_{1}=K14a18636$, in the notation from Snappy. We then show that the knot $K_{2}=K15n81556$ has unknotting number (at most) 2, and has a diagram for which one crossing change yields a diagram for $K_{1}$. Consequently, $L=K\#\overline{K}$ has a diagram for which 2 crossing changes can take us to a diagram of $K_{1}$; an isotopy then takes us to another diagram for $K_{1}$ that admits a crossing change to a diagram of $K_{2}$, which has unknotting number at most $2$. So it takes at most $2+1+2=5$ crossing changes, with isotopies, to transform our diagram of $K\#\overline{K}$ to the unknot. This unknotting sequence therefore has length $5$ (that is, $6$ knots and $5$ crossing changes); consequently, $u(K\#\overline{K})\leq 5$.
The initial diagram for $L$ in our unknotting sequence was found as the closure of a braid, using the program Snappy [9]. For our example, the particular braid ultimately produced, which yields the diagram for the first half of our unknotting sequence, is given by the braid word:
$[\underline{1},\underline{-4},2,3,3,3,2,3,2,2,4,-3,-3,-3,-3,-1,-3,-2,-3,-3]$
which represents a knot diagram with 20 crossings. Here we use Snappy’s convention that standard braid generators are listed by their (signed) subscripts; a positive entry means a generator, negative means its inverse. This braid closure is given in Figure 1; the braid is read left to right, with braid generator $1$ at bottom and $4$ at top.
Figure 1. The initial braid closure
The reader can verify in Snappy/Sage [9],[27] that this is a braid representative for $L$ by using the command K=Link(braid_closure=brd) or K=snappy.Link(braid_closure=brd) (calling the above tuple ‘brd’) and employing the simplify() and deconnect_sum() commands to recover the summands $7_{1}$ and $\overline{7_{1}}$ Alternatively, one can start from a diagram for $L=7_{1}\#\overline{7_{1}}$, on the left side of Figure 2, and arrive at the diagram of the knot built by Snappy from the braid word after a few Reidemeister moves, as indicated in Figure 2.
Figure 2. Diagrams for $7_{1}\#\overline{7_{1}}$
Crossing changes on this diagram can be carried out by changing the signs of entries in the tuple, to create a new braid word representing the knot after crossing change. We have, in the above braid word, underlined the entries relevant to our construction. Recalling that Snappy/python starts labeling entries in a tuple at ‘0’, changing the crossings at indices 0 and 1 yields the braid word
$[\underline{-1},\underline{4},2,3,3,3,2,3,2,2,4,-3,-3,-3,-3,-1,-3,-2,-3,-3]$
which Snappy identifies as belonging to the knot $K_{1}=K14a18636$. A diagram for this knot $K_{1}$ built from the closure of this braid represented by the braid word is given in Figure 3; the two crossing changes made to obtain it from the rightmost diagram in Figure 2 are circled. We also provide a schematic of an isotopy from this 20-crossing projection to a 14-crossing projection of $K_{1}$. Since unknotting number changes by at most one under crossing change, and we can recover $L$ from $K_{1}$ by reversing these $2$ crossing changes, we have $u(L)\leq u(K_{1})+2$
Figure 3. Diagrams for $K14a18636$
We now demonstrate that $K_{1}$ has unknotting number at most $3$. We do this by making use of a pair of 15-crossing diagrams of the knot $K_{2}=K15n81556$. For this discussion we utilize Dowker-Thistlethwaite (DT) codes [10] for the knots, where, again, each entry of the code representents a crossing in the diagram, and a crossing change can be achieved by changing the sign of an entry of the code. The knot $K_{2}$ has, from the authors’ earlier work on the Bernhard-Jablan conjecture [6], 100 distinct 15-crossing projections. For one of them, with DT code
$[\underline{-4},-16,24,26,18,20,28,22,-2,10,12,30,6,8,14]$,
a single crossing change, at index 0, yields the DT code
$[\underline{4},-16,24,26,18,20,28,22,-2,10,12,30,6,8,14]$,
which Snappy identifies as belonging to the knot $K_{1}=K14a18636$. Figure 4 shows how the last diagram in Figure 3, for $K14a18636$, can be transformed via Reidemeister moves and a single crossing change into the diagram for $K15n81556$ corresponding to this DT code.
Figure 4. $K14a18636$ to $K15n81556$
On the other hand, a different diagram for $K_{2}$, with DT code
$[4,12,-24,14,18,2,\underline{20},26,8,10,-28,-30,16,-6,-22]$,
when we change a single crossing, at index 6, yields the code
$[4,12,-24,14,18,2,\underline{-20},26,8,10,-28,-30,16,-6,-22]$,
which Snappy identifies as belonging to the knot $K_{3}=K12n412$. The knot $K_{3}$ has unknotting number 1 (as noted, for example, on the Knotinfo site [19]); and in fact changing the crossing in this DT code, at index 13, yields the code
$[4,12,-24,14,18,2,\underline{-20},26,8,10,-28,-30,16,\underline{6},-22]$,
which is a diagram of the unknot. A sequence of Reidemeister moves from the diagram for $K15n81556$ to the unknot, after making these two crossing changes, is shown in Figure 5. Consequently, we can obtain the unknot starting from $K_{1}=K14a18636$ with three crossing changes; first to $K_{2}=K15n81556$, then to $K_{3}=K12n412$, and finally to the unknot.
Figure 5. $K15n81556$ to the unknot
This establishes that $u(K_{1})=u(K14a18636)\leq 3$. Therefore,
$u(7_{1}\#\overline{7_{1}})=u(L)\leq u(K_{1})+2\leq 3+2=5$
Since $u(7_{1})+u(\overline{7_{1}})=3+3=6$, this shows that $u(7_{1}\#\overline{7_{1}})<u(7_{1})+u(\overline{7_{1}})$, establishing the failure of additivity for these knots. This completes the proof of Theorem 1.2. $\square$
Extending our main example to an infinite class of examples takes little extra work. A knot $K_{1}$ is called Gordian adjacent to the knot $K_{2}$ [11] if $u(K_{2})>u(K_{1})$ and there are $u(K_{2})-u(K_{1})$ crossing changes on a diagram for $K_{2}$ that results in a diagram for $K_{1}$. Put differently, $K_{2}$ possesses a minimal unknotting sequence which contains the knot $K_{1}$. For any pair of knots $L_{1},L_{2}$ that the knot $7_{1}$ is Gordian adjacent to, we can build an unknotting sequence for $L_{1}\#\overline{L_{2}}$ using fewer than $u(L_{1})+u(L_{2})$ crossing changes. Starting with $L_{1}\#\overline{L_{2}}$, applying $u(L_{1})-3=u(L_{1})-u(7_{1})$ crossing changes to the $L_{1}$-summand, and $u(L_{2})-3$ crossing changes to the $\overline{L_{2}}$ summand will yield a diagram for $7_{1}\#\overline{7_{1}}$, and this knot can be unknotted with $5$ crossing changes (after isotopy). Conequently, $(u(L_{1})-3)+(u(L_{2})-3)+5=u(L_{1})+u(L_{2})-1$ crossing changes suffice to unknot $L_{1}\#\overline{L_{2}}$, and so $u(L_{1}\#\overline{L_{2}})\leq u(L_{1})+u(L_{2})-1<u(L_{1})+u(L_{2})$.
One family of knots satisfying this condition is the set of torus knots $T(2,2k+1)$ with $k\geq 3$. These knots have unknotting number $(2-1)((2k+1)-1)/2=k$, and changing any $k-3$ crossings in their standard diagrams will yield $T(2,7)=7_{1}$ (since changing a crossing in the standard diagram for $T(2,N)$ will, after a Reidemeister II move, yield the standard diagram for $T(2,N-2)$, and we are then done by induction). This yields Corollary 1.3. $\square$
Even more, Feller [11] shows that for the torus knots $T(a,b)$ with $gcd(a,b)=1$ and $a\geq 2$, $b\geq 7$ the knot $T(2,7)=7_{1}$ is Gordian adjacent to each of these knots, and, in addition, $T(2,7)$ is Gordian adjacent to the knot $T(3,5)$. All of these knots can therefore also be used to build a connected sum that fails additivity.
Corollary 2.1.
For all integers $a_{1},a_{2}\geq 2$ and $b_{1},b_{2}\geq 7$ (or $b_{i}\geq 5$ if $a_{i}=3$) with $\gcd(a_{i},b_{i})=1$ for $i=1,2$, we have
$u(T(a_{1},b_{1})\#\overline{T(a_{2},b_{2})})<u(T(a_{1},b_{1}))+u(T(a_{2},b_{2}))\ .$
For an example which is not a torus knot, we can use the knot $10_{139}$ (which, in fact, was a summand of the first example we found in proving Theorem 1.2; see Section 3 below). The knot $L_{1}=10_{139}$ is a non-alternating hyperbolic knot; it was shown by Kawamura [15] that $u(L_{1})=4$. In fact, Kawamura showed, even more, that $L_{1}$ has smooth 4-genus $g_{4}(L_{1})=4$ (and $g_{4}(L)\leq u(L)$ for every knot $L$ [21]); the 4-genus of $L_{1}$ was also established by different means (using their tau invariant) by Ozsváth and Szabó [24]. One DT code for $10_{139}$ is given by
[12, 14, -10, -20, -16, 18, 2, -8, 4, -6],
and changing the sign of the entry at index 4 yields the DT code
[12, 14, -10, -20, 16, 18, 2, -8, 4, -6],
which belongs to the knot $7_{1}$. As a result, an unknotting sequence for $7_{1}$ (with $3$ crossing changes) can be augmented with this crossing change to yield an unknotting sequence for $10_{139}$ with $4$ crossing changes, that is, a minimal unknotting sequence. Consequently, $10_{139}$ can play the role of $L_{1}$ (and $L_{2}$) in the above discussion, yielding:
Corollary 2.2.
$u(7_{1}\#\overline{10_{139}})<u(7_{1})+u(10_{139})$ and $u(10_{139}\#\overline{10_{139}})<u(10_{139})+u(10_{139})$
We expect that there are many other knots whose unknotting numbers are known and that contain $7_{1}$ in a minimal unknotting sequence, which would then yield other connected sums that fail additivity. For example, any intermediate knot $K$ in a Gordian adjacency from the torus knot $T(p,q)$ to the knot $7_{1}$ also has $7_{1}$ Gordian adjacent to it, and so can be used as a summand in such examples. Several authors have explored the construction of minimal unknotting sequences for torus knots [3],[33], and the problem of Gordian adjacency for torus knots [11],[5]. Feller’s work [11] shows that all torus knots except (possibly) $T(2,3),T(2,5),T(3,4),T(4,5)$ and $T(5,6)$ have $7_{1}=T(2,7)$ Gordian adjacent to them; $T(2,3),T(2,5)$ and $T(3,4)$ cannot, because their unknotting numbers are too low. The cases $T(4,5)$ and $T(5,6)$ are unresolved.
3. How (and why) the main example was found
The process of finding the diagram for $7_{1}\#\overline{7_{1}}$ and the discovery of its unexpectedly low unknotting number had its genesis in another ongoing project of the authors: our search for an example to answer a different question, namely whether or not unknotting number is invariant under mutation. This is Problem 1.69(C) in Kirby’s problem list [16]. Our approach to that problem included building a collection of knots which would have unknotting number 4 (if unknotting number were additive), but which would have comparatively small signature. While the inequality $|\sigma(K)|/2\leq u(K)$ [21] is useful in bounding unknotting number from below, it can be difficult to find knots where the difference between these two quantities is known to be greater than 1 (which we needed for our search), although work of Stoimenow [34] is a significant outlier, which deserves to be more widely known. On the other hand, the facts that signature is additive under connected sum, and that the signature of the mirror image of a knot $K$ is the negative of the signature of $K$ (see, e.g., [21]), means that is it a routine matter to build connected sums where the unknotting number would be large, if unknotting number were additive, but where the signature is small. It was for this reason that we began to search for knots a few crossing changes away from connected sums of low crossing number knots with known unknotting number, in order to build a collection of new knots with (if unknotting number were additive) high enough unknotting number to be useful for our search.
Our approach to finding the new knots was to adopt the process we had followed for two previous projects, where we settled the Bernhard-Jablan Conjecture (in the negative) [6] and where we completed the computation of the smooth 4-genus of the prime knots through 12 crossings [7]. We started with pairs of knots, one with known unknotting number 3 and the other with known unknotting number 4, took their connected sum, and then built diagrams for these sums for which crossing changes stood a good chance of building prime knots, with low enough (after simplification) crossing number that Snappy could recognize them (that is, crossing number 15 or less). We chose unknotting numbers 3 and 4 so that we were starting with knots that were likely to have unknotting number 7, and then randomly changed three crossings so that the resulting knots ought to have unknotting number (at least) 4. We chose to build the diagrams by starting with the standard diagrams for the connected sums, randomly performing some crossing-number-increasing Reidemeister moves using Snappy’s backtrack() command, and then converting them to braids using the braid_word() command. For this last step, Snappy uses Vogel’s algorithm [36] to convert the knot diagram into a closed braid. We chose this approach because (a) Snappy can quickly carry out both of these operations and (b) our experience with the previous two projects had shown us that working with braid representatives built in this way succeeds in building diagrams that are randomized enough to yield good results (that is, many different knot types) when we change crossings in the braid representations. Experience also showed that by changing 3 crossings, when starting from a connected sum in this way, resulted in a good proportion of the knots that were generated being prime; 2 changes was typically insufficient. Using braid words, changing crossings is very quick; it amounts to changing the signs of entries in the braid representation.
Our work in [6] had provided computations of upper bounds on the unknotting number of all of the prime knots through 15 crossings (what we called, in that paper, their weak BJ-unknotting number, that is, their unknotting number if the Bernhard-Jablan conjecture (and additivity of unknotting number!) were true). We could have high confidence that the vast majority of these computations represented the actual unknotting number of the knots, based on comparison with known values for lower crossing number knots. We compared the knots coming from the randomized procedure to our list of ‘weak BJ-unknotting number 4’ knots to create a group of what we called ‘conditionally unknotting number four’ knots. Working with several million braids yielded 240 such 14- and 15-crossing knots. One of these was the knot $K14a18636$.
We then began to reverse the search, using these 240 knots as initial knots to carry out crossing changes on their braid representatives, in the same manner in which the connected sum diagrams had been built. This is where things got interesting! For the mutation project, our goal was to show that one of a collection of target knots had unknotting number 3, by showing that it was crossing-adjacent to a knot with unknotting number 4. Prior such searches using knots with unknotting number known to be 4 had failed to make such a connection, and so our plan was to search among knots one crossing change away from these 240 knots with conditional unknotting number 4, instead. Connecting one of the 240 initial knots with one of these mutant ‘targets’ would have proved that either unknotting number is not mutation invariant or that the knot among our 240 that made the connection could not actually have unknotting number 4, and so, as argued above, unknotting number would not be additive under connected sum. (But, were such a thing to happen, we would not know which of the two possibilities was true!)
At the start of this phase of the project it was not at all clear what it would produce. The most likely prospect seemed to be that we would never find a connection between our 240 initial knots and our list of target knots, if only because both conjectures might be true. Our expectation, if the project succeeded, was that it would be because the target knot we connected with had unknotting number 3, and so unknotting number would not be mutation invariant. It came as quite a surprise then when, instead, this reverse search found a crossing change from one of the 240 knots we had identified, namely $K14a18636$, to the knot $K15n81556$, which had a known, and unexpectedly low, weak BJ-unknotting number of 2.
The first example we found started from a closed braid representation of the knot $7_{1}\#10_{139}$; three crossing changes, that our search had found, transformed this to a representative of the knot $K14a18636$. One of those three crossing changes transformed the braid into a braid representative for the knot $7_{1}\#\overline{7_{1}}$, and so this braid word contained two crossing changes to a representative of the knot $K14a18636$. Initially, the braid word had length 113, and so represented a 113-crossing diagram. But several simplifications of the braid word could, it turned out, be carried out. These mostly involved Reidmeister I moves ‘hidden’ by a sequence of Reidemeister II moves. Their discovery was aided by the program KnotPlot [32], to visualize long stretches of the braid. This enabled the authors to greatly reduce the number of crossings needed (from 113 to 20), and the number of strands in the braid needed (from 14 to 5). This resulted in the diagram and braid word used in our discussion in Section 2.
This example, of course, destroyed the hypothesis underlying our mutation project; but that would seem to be a small price to pay, for the result that it achieved.
4. The road forward
With the additivity of unknotting number being settled in the negative, it might seem like the quest to build a more comprehensive understanding of unknotting number may have hit a bump in the road. Knowing for sure that unknotting number does not behave well with respect to connected sum means that the chances for a purely computational approach to determining unknotting numbers, as one could argue the other invariants mentioned in the introduction enjoy, seems even further out of reach. Further, we cannot rely on knowledge of unknotting numbers for prime knots to provide all of the information we need to compute the invariant more generally.
As the authors of the reinforcment learning approach [2] had postulated, additivity under connected sum presented a potentially powerful opportunity to compute the unknotting numbers of knots with relatively small crossing number but unknown unknotting numbers, by building ‘test’ knots, as connected sums, which effectively decouple the unknotting number from the signature (and, as it happens, the 4-genus), to change crossings in. With the loss of this technique, we need new tools to determine the unknotting numbers, for example, of the nine 10-crossing prime knots whose values we still do not know.
On the other hand, knowing that additivity of unknotting number can in fact fail makes it more interesting to discover when it does not fail. That is, establishing that $u(K\#K^{\prime})=u(K)+u(K^{\prime})$ for families of knots has increased importance, since we can no longer expect this to hold in general. There is also the question of how badly additivity fails; is it, perhaps, even possible that the unknotting number of a connected sum can be lower than that of its summand?
Question 4.1.
Does $u(K\#K^{\prime})\geq\max\{u(K),u(K^{\prime})\}$ for all knots $K$ and $K^{\prime}$
The fact that we could bootstrap an initial, single, example of non-additivity to find infinitely may knots $K$ for which $u(K\#7_{1})<u(K)+u(7_{1})$ raises the question of whether or not every knot $K$ can be paired with a ‘poison’ knot $K^{\prime}$ so that $u(K\#K^{\prime})<u(K)+u(K^{\prime})$. So we ask:
Question 4.2.
Does there exist a non-trivial knot $K$ for which $u(K\#K^{\prime})=u(K)+u(K^{\prime})$ for every knot $K^{\prime}$
While we now know that $u(7_{1}\#\overline{7_{1}})<u(7_{1})+u(\overline{7_{1}})$, and we can build infinitely many examples of this phenomenon from this, all of our examples currently rely on a sequence of $5$ crossing changes that unknot $7_{1}\#\overline{7_{1}}$ We would certainly like to understand why this sequence occurred! What, if anything, is special about the knot $7_{1}$ that might make this happen? The intermediate knot $K_{1}=K14a18636$, which played a central role in establishing the non-additivity of unknotting number, has, from our work in [6], a weak BJ-unknotting number of 4, yet the unknotting sequence above shows that it has unknotting number at most 3. So the knot $K_{1}$ is, in fact, a counterexample to the weak Bernhard-Jablan conjecture [6]. That is, no crossing change in its (unique) minimal crossing diagram lowers the weak BJ-unknotting number. This may point the way to building more examples, by aiming to find a BJ-unknotting counterexample in an unknotting sequence. We note that there are (at least) six other knots that the authors have identified with weak BJ-unknotting 4 and unknotting number $\leq 3$: they are the knots $K13n1669$, $K14a2644$, $K14a11300$, $K14a11565$, $K14a12383$, and $K14a13809$. Each is crossing-adjacent to a 15-crossing diagram for a knot $K$ with $u(K)\leq 2$ (the corresponding unnotting number (at most) 2 knots are, respectively, $K15n9318$, $K15n9318$, $K15n35674$, $K15n82797$, $K15n4866$, and $K15n129095$), and so each of these knots has unknotting number at most $3$. These could, like $K14a18636$, be useful targets for building other unexpectedly short unknotting sequences for connected sums.
Finally, we are left with the question:
Question 4.3.
What is the actual unknotting number of $7_{1}\#\overline{7_{1}}$
We now know that the unknotting number of $7_{1}\#\overline{7_{1}}$ is at most $5$, and, by Scharlemann’s result [28], it is at least 2. Could it actually be lower than $5$?
5. Verification code
For the convenience of the reader, we give the code (written for Sage; remove the occurances of ‘snappy’ to run in Snappy) which will verify all of the identifications made for our example. We also include the output from this code, as carried out in Sage. The reader’s output may vary slightly, due to some randomness built into some of Snappy’s routines. Note: When cutting and pasting from the code below, the quotation marks may need to be retyped by the reader before pasting into Snappy/Sage.
import snappy
brd=[1,-4,2,3,3,3,2,3,2,2,4,-3,-3,-3,-3,-1,-3,-2,-3,-3]
print(’First braid word (for L) is’)
print(str(brd))
L=snappy.Link(braid_closure=brd)  ## building the connected sum
L.simplify(’global’)
print(’Summands of L, and their number of crossings:’)
PP=K.deconnect_sum()       ## this yields 7_1 and mirror(7_1)
PP
print(’Attempting to identify summands:’)
PP[0].exterior().identify()   ## these may be "[]"; Snappy does
PP[1].exterior().identify()   ## not always succeed in ’recognizing’
                              ## non-hyperbolic knots.
print(’First summand knot group is ’)
PP[0].exterior().fundamental_group()
print(’Second summand knot group is ’)
PP[1].exterior().fundamental_group()
## both of these will yield <a,b|aaaaaaabb> (or <a,b|aabbbbbbb>);
## 7_1 is the only knot with this group.
brd2=brd[:]    ## implementing the 2 crossing changes
brd2[0]=-brd2[0]
brd2[1]=-brd2[1]
LA=snappy.Link(braid_closure=brd2)  ## this is K14a18636
print(’Second braid word (for LA) is ’)
print(str(brd2))
LA.simplify(’global’)
print(’LA is ’+str(LA.exterior().identify()))
DTC=[4,-16,24,26,18,20,28,22,-2,10,12,30,6,8,14]
DTCB=DTC[:]
print(’DT code for KA is ’)
print(str(DTC))
DTCB[0]=-DTCB[0]   ## change one crossing
print(’DT code for KB is ’)
print(str(DTCB))
KA=snappy.Link(’DT:’+str(DTC))     ## this is K14a18636
KB=snappy.Link(’DT:’+str(DTCB))    ## this is K15n81556
print(’KA is ’+str(KA.exterior().identify()))
print(’KB is ’+str(KB.exterior().identify()))
print(’Snappy check: LA and KA are the same knot? ’
+str(LA.exterior().is_isometric_to(KA.exterior())))
DTD=[4,12,-24,14,18,2,20,26,8,10,-28,-30,16,-6,-22]
print(’DT code for KC is ’)
print(str(DTD))
DTDB=DTD[:]
DTDB[6]=-DTDB[6]   ## change one crossing
print(’DT code for KD is ’)
print(str(DTDB))
KC=snappy.Link(’DT:’+str(DTD))     ## this is K15n81556
KD=snappy.Link(’DT:’+str(DTDB))    ## this is K12n412
print(’KC is ’+str(KC.exterior().identify()))
print(’KD is ’+str(KD.exterior().identify()))
print(’Snappy check: KB and KC are the same knot? ’
+str(KB.exterior().is_isometric_to(KC.exterior())))
DTE=DTDB[:]
DTE[13]=-DTE[13]
print(’DT code for KE is ’)
print(str(DTE))
KE=snappy.Link(’DT:’+str(DTE))    ## this is the unknot
print(’KE is ’+str(KE.exterior().identify()))
print(’Knot group for KE:’)
KE.exterior().fundamental_group()
## only the unknot has cyclic knot group
Sample output from this code:
>  First braid word (for L) is
>  [1, -4, 2, 3, 3, 3, 2, 3, 2, 2, 4, -3, -3, -3, -3, -1, -3, -2, -3, -3]
>  True
>  Summands of L, and their number of crossings:
>  [<Link: 1 comp; 7 cross>, <Link: 1 comp; 7 cross>]
>  Attempting to identify summands:
>  [7_1(0,0), K7a7(0,0)]
>  [7_1(0,0), K7a7(0,0)]
>  First summand knot group is
>  Generators:
>     a,b
>  Relators:
>     aaaaaaabb
>  Second summand knot group is
>  Generators:
>     a,b
>  Relators:
>     aabbbbbbb
>  Second braid word (for LA) is
>  [-1, 4, 2, 3, 3, 3, 2, 3, 2, 2, 4, -3, -3, -3, -3, -1, -3, -2, -3, -3]
>  True
>  LA is [K14a18636(0,0)]
>  DT code for KA is
>  [4, -16, 24, 26, 18, 20, 28, 22, -2, 10, 12, 30, 6, 8, 14]
>  DT code for KB is
>  [-4, -16, 24, 26, 18, 20, 28, 22, -2, 10, 12, 30, 6, 8, 14]
>  KA is [K14a18636(0,0)]
>  KB is [K15n81556(0,0)]
>  Snappy check: LA and KA are the same knot? True
>  DT code for KC is
>  [4, 12, -24, 14, 18, 2, 20, 26, 8, 10, -28, -30, 16, -6, -22]
>  DT code for KD is
>  [4, 12, -24, 14, 18, 2, -20, 26, 8, 10, -28, -30, 16, -6, -22]
>  KC is [K15n81556(0,0)]
>  KD is [K12n412(0,0)]
>  Snappy check: KB and KC are the same knot? True
>  DT code for KE is
>  [4, 12, -24, 14, 18, 2, -20, 26, 8, 10, -28, -30, 16, 6, -22]
>  KE is []
>  Knot group for KE:
>  Generators:
>     a
>  Relators:
6. Acknowledgements
The first author acknowledges support by a grant from the Simons Foundation (Collaboration Grant number 525802). The second author acknowledges support by a grant from the Simons Foundation (Collaboration Grant number 581433). The authors also acknowledge the support of the Holland Computing Center at the University of Nebraska, which provided computing facilities on which the some of the work toward this project was carried out.